OpenStudy (anonymous):

Sam and Chris went to “Lots O Fun” to play laser tag and video games for Chris’s birthday. Sam played 3 games of laser tag, 5 video games, and spent $17 total. Chris played 4 games of laser tag, 7 video games, and spent $23 total. How much does one video game cost to play?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

First, pick variable names for the unknowns. can you do that ?

OpenStudy (anonymous):

l for laser tag and v for video games

OpenStudy (phi):

ok, but I would not use l because it looks like one (1) . how about L ? also, be more clear: L and v are the COST of each type of game: L= cost of laser tag games v = cost of video games next, look at this sentence: Sam played 3 games of laser tag, 5 video games, and spent $17 total. can you make an equation (showing cost of games plus cost of games = total cost) ?

OpenStudy (phi):

if one laser tag game costs L dollars, how much does 3 games cost ?

OpenStudy (anonymous):

3L?

OpenStudy (phi):

yes. Sam played 3 games of laser tag, and it costs 3L (he paid L dollars for each game... 3 times he dug into his pocket and paid for a laser tag game... paying L dollars each time: L+L+L or 3L dollars

OpenStudy (phi):

Sam played 3 games of laser tag, 5 video games, and spent $17 total. can you make an equation (showing cost of games plus cost of games = total cost) ?

OpenStudy (phi):

if 1 video game costs v dollars, how much does 5 video games cost ?

OpenStudy (anonymous):

5v. So would it be 3L+5v=17?

OpenStudy (phi):

yes. you now have 1 equation and 2 unknowns It is an important fact that you need as many equations as unknowns if you want to solve the system of equations. you have 2 unknowns: L and v (cost of laser tag and cost of video) you need one more equation. use Chris played 4 games of laser tag, 7 video games, and spent $23 total.

OpenStudy (anonymous):

4L+7v=23

OpenStudy (phi):

yes. now we are back in the land of algebra: 3L+5v=17 4L+7v=23 these are a bit ugly to solve but do-able. How would you solve them ?

OpenStudy (phi):

are you going to use substitution or elimination ?

OpenStudy (anonymous):

move 5v to the other side and substitute 17-5v for y in the second equation and solve for L

OpenStudy (phi):

move 5v to the other side and substitute 17-5v for y <grumble, grumble> there is no "y". Remember, be *careful* 3L+5v=17 4L+7v=23 using the 1st equation, move 5v to the other side is right. you get: 3L = 17 - 5v this says 3L is equal to 17 -5v what is 1 L equal to? divide both sides by 3 to get 1L 3L/3 = (17-5v)/3 (notice we divide the whole side by 3... so we use parens) L = (17-5v)/3 now substitute this in the 2nd equation. can you do this ?

OpenStudy (anonymous):

which wad the 2nd equation?

OpenStudy (phi):

3L+5v=17 4L+7v=23

OpenStudy (phi):

solve for L in the first equation: 3L+5v=17 you get L = (17-5v)/3 substitute for L in the 2nd equation 4L+7v=23 what do you get ?

OpenStudy (anonymous):

I got 15/13?

OpenStudy (phi):

let's do it step by step. after substituting L = (17-5v)/3 into 4L+7v=23 what do you get ?

OpenStudy (phi):

the idea is not think too hard (make each step simple) so "erase" the L and put in (17-5v)/3 in 4L+7v=23 you get \[ 4 \left( \frac{(17-5v)}{3} \right)+ 7v =23 \]

OpenStudy (phi):

I don't like factions (I assume you don't either). One way to get rid of the 3 in the bottom is to multiply the equation (both sides, *all* terms) by 3 like this \[ 4 \cdot 3 \left( \frac{(17-5v)}{3} \right)+ 3 \cdot 7v =23 \cdot 3 \]

OpenStudy (phi):

the 3/3 "cancels" in the first term, so you get \[ 4 \cdot \cancel{3} \cdot \frac{(17-5v)}{\cancel{3} }+ 3 \cdot 7v =23 \cdot 3\\ 4 (17-5v) + 21v = 69 \]

OpenStudy (phi):

solve for v first, distribute the 4, then combine "like terms"